OpenStudy (katiethompsonnn):

Simplify.. 2(2/2-x)[-2/(2-x)^2]

OpenStudy (katiethompsonnn):

@Hero please help

hero (hero):

\[2\left(\dfrac{2}{2-x}\right)\left[\dfrac{-2}{(2-x)^2}\right]\]

hero (hero):

That's what it looks like you posted.

OpenStudy (katiethompsonnn):

yeah thats what my problem looks like

hero (hero):

Looks like all you have to do is multiply

OpenStudy (pythagoras123):

For multiplying of algebraic fractions: \[\frac{ a }{ b } \times \frac{ c }{ d }=\frac{ ac }{ bd }\]